### v2

# Dynamical de Sitter conjecture and quintessence

Abstract
The de Sitter conjecture yields a severe bound on possible vacua for a consistent quantum gravity. We extend the de Sitter conjecture by taking into account dynamics of the scalar field. We then apply such an extended de Sitter conjecture to a quintessence model of inflation for which dynamics of the scalar field is essential, and obtain an allowed region of parameters of the scalar potential wider than previously considered cases with the conventional de Sitter conjecture. The new bounds in the swampland conjecture could have implications in several situations to construct compactification models.
pacs:
04.50.-h, 11.25.Mj, 98.80.Cq, 95.36.+x
I Introduction
String theory is one of the candidates for quantum field theories consistent with theory of gravity. Since it is defined in higher-dimensional spacetimes, a compactification process becomes necessary to ensure compatibility with our four-dimensional universe. Each compactification represents a different low-energy effective field theory in four dimensions, which generates what is called the “landscape” of string theory Susskind:2003kw ; Banks:2003es ; Denef:2004ze ; Kallosh:2004yh ; Vafa:2005ui ; Denef:2008wq ; Brennan:2017rbf ; vanBeest:2021lhn ; Grana:2021zvf ; Agmon:2022thq . Although the four-dimensional quantum field theory describing the standard model of particle physics has succeeded in explaining various interactions, it is also true that numerous problems remain in field theories beyond the standard model, such as the need for parameter fine-tuning to determine coupling constants and particle masses.
In recent developments of string theory, it has been suggested that not all quantum field theories are necessarily consistent with gravitational theories, and there exist conditions for four-dimensional field theories to be gravitationally consistent. Field theories that do not satisfy these conditions are referred to as residing in the “swampland” Vafa:2005ui ; Ooguri:2006in ; Ooguri:2016pdq ; Brennan:2017rbf ; Obied:2018sgi ; Agrawal:2018own ; Ooguri:2018wrx ; Andriot:2018wzk ; Palti:2019pca ; Bedroya:2019snp ; Agmon:2022thq . Some solutions to higher-dimensional string theories become theories within the landscape through compactification, while others become theories residing outside the landscape, i.e., in the swampland. The latter are no longer consistent with gravitational theories. The classification of quantum field theories consistent with quantum gravity potentially provides powerful theoretical tools, offering strategies to resolve fine-tuning problems Heckman:2019bzm and enabling predictions about cosmological observations by clarifying the swampland region, such as the dynamics of early universe inflation and the nature of dark energy Agrawal:2019dlm ; Bedroya:2019snp ; Vafa:2019evj ; McNamara:2020uza ; vanBeest:2021lhn ; Agmon:2022thq . Although string theory might provide the field theory of the standard model of particle physics, no one knows the full extent of the landscape. If the landscape were to encompass all quantum field theories, string theory would find it challenging to provide insights into field theories beyond the standard model. However, we cannot obtain the low-energy limits of all string-theory compactifications, and it is becoming increasingly clear that numerous field theories exist outside the landscape Vafa:2005ui ; Palti:2019pca .
Consequently, the most critical issue in current string theory is determining the boundary between the landscape and the swampland. However, given the countless possible compactifications of string theory, definitively establishing this boundary is not straightforward Brennan:2017rbf . Therefore, there are systematic investigations of reliable compactifications that provide realistic four-dimensional universe models and standard model consistent with observations and experiments, empirically confirming the characteristics of four-dimensional field theories they share. From these data, studies are developing to hypothesize criteria for determining which field theories belong to the landscape or swampland and to define the boundary between them Brennan:2017rbf ; Palti:2019pca ; Bedroya:2019snp .
As an example of determining such a boundary, the so-called de Sitter conjecture has been proposed.This conjecture addresses the possibility that the de Sitter vacuum becomes part of the swampland in the quantum field theory under consideration, and it is derived from the field equations and energy conditions of string theory Obied:2018sgi . The de Sitter conjecture is further supported by discussions of distance conjectures related to the de Sitter spacetime entropy and the dynamics of extra dimensions Ooguri:2018wrx . For instance, as extra dimensions increase in size, light particles emerge on the four-dimensional spacetime Ooguri:2006in , and their contributions impact the dynamics of early universe inflation. These new particles are colloquially referred to as the “tower of light states” Ooguri:2006in and play a role in constraining inflationary models Agrawal:2018own ; Garg:2018reu . The de Sitter conjecture provides constraints on the low-energy effective theories of consistent quantum gravitational theories, and these constraints (that is, the criteria for determining whether a theory belongs to the landscape or swampland) stem from the current accelerating expansion of the universe. In other words, the low-energy effective theories derived from consistent quantum gravity theories are evaluated on the basis of whether they can reproduce the observed accelerating universe and the inflation of the early universe without contradiction. The criteria determine whether a quantum field theory resides in the landscape or in the swampland. Naturally, if a low-energy effective theory realizes a meta-stable de Sitter vacuum Kachru:2003aw ; Kachru:2003sx and accelerated expansion at the present universe, it would be considered part of the landscape. Conversely, if it fails to do so, the four-dimensional quantum field theory would be deemed to reside in the swampland Obied:2018sgi ; Andriot:2018wzk ; Garg:2018reu ; Palti:2019pca .
The de Sitter conjecture gives a lower bound on the gradient of potentials in positive regions Obied:2018sgi . The scalar potential of a theory coupled to gravity have to satisfy a bound on its derivatives with respect to scalar fields
$|\nabla V|>c\,V\,,$ (1)
where $c>0$ is of order 1 in Planck units. The number $c$ generally depends on the number of dimensions of background space-time and internal space. This bound on $\left|\nabla V\right|/V$ for compactifications of string theory is in fact a consequence of strong energy condition (SEC) in higher dimensions and field equations. The criteria can be derived by a simple generalization of Maldacena & Nuñez Maldacena:2000mw and Steinhardt-Wesley Steinhardt:2008nk no-go theorems that one can place a bound on the gradient of the potential.
As mentioned above, the de Sitter conjecture is fundamentally constructed based on the field equations of the original higher-dimensional theory, the low-energy effective theory obtained after compactification, and the SEC Obied:2018sgi . In this process, the time derivative of the scalar field is set to zero. The scalar field represents the moduli in the higher-dimensional theory, corresponding to the scale of the internal space, is fixed Obied:2018sgi . If the scalar field coming from moduli triggers an inflaton, the condition of dropping off the time derivative would be consistent with the slow-roll condition of inflation.
The early universe inflationary expansion are achieved by the dynamics of scalar fields, particularly through scalar fields which possess minimum or local minimum in their potential Kachru:2003aw ; Kachru:2003sx . The potential energy at these minima or local minima of scalar fields is equivalent to a positive cosmological constant in four-dimensional effective theories, thus leading to stable or meta-stable de Sitter spacetimes. In string theory, there are numerous efforts to make stable or meta-stable de Sitter vacua Kachru:2003aw ; Kachru:2003sx , but an overwhelming number of problems still have emerged. Within the issues of reproducing stable or meta-stable de Sitter vacua in string theory, recent proposals have focused on constraints imposed on scalar field potentials to avoid residing in the swampland.
On the other hand, the current accelerating cosmic expansion should be explained by a potential of the scalar field that lacks minimum or local minimum; If the potential remains consistently positive and the contribution from its kinetic term is not overly significant, one can still explain an accelerating universe consistent with current observational data. Such scalar fields are often referred to as quintessence Caldwell:1997ii ; Carroll:1998zi ; Zlatev:1998tr ; Chiba:1999ka ; Tsujikawa:2013fta . The simplest and most natural choice of the potential for quintessence models is given by an exponential function Barreiro:1999zs ; Raveri:2018ddi
$V=V_{0}\,\exp\left(-\lambda\tau\right)\,,$ (2)
where $\tau$ is a scalar field and $\lambda$ is a constant slope. However, a puzzle issue has recently emerged with the quintessence models with the above scalar potential. The allowed region $\lambda<0.60$ of the potential slope $\lambda$ in the numerical solutions of the field equations in four-dimensional dynamical systems Andriot:2024jsh ; Rudelius:2022gbz ; Shiu:2023nph ; Copeland:1997et ; Ferreira:1997hj ; Bahamonde:2017ize ; Alestas:2024gxe ; Casas:2024oak suggested by observational facts of supernovae DES:2024jxu and BAO DESI:2024mwx is not consistent with the result $\lambda\geq\sqrt{2}$ constrained bythe de Sitter conjecture Obied:2018sgi . However, when considering quintessence, vanishing the time derivative of the scalar field is not necessarily a condition that must be satisfied. Here there is also a problem with quintessence as dark energy/late time accelerated expansion is that it is guaranteed to make $H_{0}$ tension. These issues are explored in Colgain:2019joh ; Banerjee:2020xcn ; Lee:2022cyh . They slowly generalized the statement from the potential (2), to generic potentials, and beyond quintessence to the Horndeski class of models.
There are some comments on the contents other than those mentioned above. For a multi-field scenario, the late-time attractor is necessarily the scaling solution Collinucci:2004iw such that the scale factor grows like a power of cosmological time Shiu:2023fhb ; Shiu:2024sbe . When the lower-dimensional dilaton is one of the rolling fields then the strong de Sitter conjecture is always fulfilled in the presence of just positive potentials Shiu:2023fhb ; Shiu:2023nph . There is also simple criteria even if we have negative potential terms Shiu:2023nph ; VanRiet:2023cca . One can also has some constraints about contracting universes with negative-definite potentials Shiu:2023yzt ; Uzawa:2018sal .
This note studies the de Sitter conjecture with dynamical scalar field, when removing the assumption of neglecting the time derivative on it. When the time derivative of scalar field is non-zero, the energy conditions and field theory equations of higher-dimensional and the low energy effective theories after compactification cannot neglect the kinetic term of scalar field, and these contributions give a correction to the de Sitter conjecture in Eq. (1). We find that if we apply such “dynamical de Sitter conjecture” to the quintessence model in which the time derivative is not negligible, we obtain an allowed region of the parameters broader than previously considered ($\lambda\geq\sqrt{2}$). Our results provide a solution to the aforementioned discrepancy between observations and the de Sitter conjecture, owing to expanding the allowed region of the parameter space for the scalar field potential.
In section II, we revisit the original de Sitter conjecture, explicitly presenting the equations of motion and strong energy conditions without neglecting the time derivative terms of the scalar field. We discuss how the kinetic terms of the scalar field contribute to the criteria of de Sitter conjecture. Section III addresses the quintessence model when the potential of the scalar field is expressed exponentially and determines the allowed region of potential slopes when applying the dynamical de Sitter conjecture to this model. In Section IV, we give cosmological considerations and future implications based on our discussions.
II de Sitter conjecture
In this section, we discuss the de Sitter conjecture or swampland criteria. After reviewing it in the first subsection, we extend it in the case of scalar field is dynamical in the second subsection.
II.1 Extension of Swampand criteria
Now we suppose a theory in $D$ dimensions and compactify it on a $(D-d)$-dimensional space with a metric $g_{mn}$ and a warp factor $\Omega$ as,
$ds^{2}=g_{MN}dx^{M}dx^{N}=\Omega^{2}(t,y)\left[-dt^{2}+a^{2}(t)\delta_{\alpha\beta}dx^{\alpha}dx^{\beta}\right]+g_{mn}(t,y)dy^{m}dy^{n}\,,$ (3)
where $g_{mn}$ denotes the metric of internal space, $x^{M}$ , $x^{\alpha}$  and $y^{m}$  are the coordinates of $D$-dimensional spacetime, $(d-1)$-dimensional space, $(D-d)$-dimensional internal space, respectively. We assume that the warp factor $\Omega$ and the internal space metric $g_{mn}$ depend only on the coordinates of the internal space $y^{m}$ and time $t$,.
We fix a residual diffeomorphism by choosing the $d$-dimensional gravitational constant $\kappa_{d}$ to be equal to $D$-dimensional one $\kappa_{D}$. This implies Obied:2018sgi
$\int d^{D-d}y\,\Omega^{d-2}\sqrt{{\rm det}\,g_{mn}}=1\,.$ (4)
Hence we have
$\frac{\partial^{2}}{\partial t^{2}}\left(\int d^{D-d}y\Omega^{d-2}\sqrt{\det g_{mn}}\right)=\int d^{D-d}y\Omega^{d-2}\sqrt{\det g_{mn}}\left[\frac{1}{2}\left(g^{mn}\ddot{g}_{mn}+\dot{g}^{mn}\dot{g}_{mn}\right)\right.$
$\left.~{}~{}~{}~{}~{}~{}+\frac{1}{4}\left(g^{mn}\dot{g}_{mn}\right)^{2}+(d-2)\Omega^{-1}\ddot{\Omega}+(d-3)\Omega^{-2}\dot{\Omega}^{2}\right]=0\,,$ (5)
where we will assume time dependences
$\dot{\Omega}\equiv\partial_{t}\Omega\neq 0\,,~{}~{}~{}~{}~{}\dot{g}_{mn}\equiv\partial_{t}g_{mn}\neq 0\,.$ (6)
The computation of the Ricci tensor $R_{tt}$ corresponding the background metric (3) gives
$R_{tt}$ $=$ $-\left(d-1\right)\left[a^{-1}\ddot{a}+\left(a\,\Omega\right)^{-1}\dot{a}\,\dot{\Omega}-\Omega^{-2}\,\dot{\Omega}^{2}+\Omega^{-1}\,\ddot{\Omega}\right]$ (7)
$-\frac{1}{2}\left(g^{mn}\ddot{g}_{mn}+\dot{g}^{mn}\dot{g}_{mn}-\Omega^{-1}\,g^{mn}\dot{g}_{mn}\,\dot{\Omega}-\frac{1}{2}g^{mm^{\prime}}\,g^{nn^{\prime}}\,\dot{g}_{nm^{\prime}}\dot{g}_{mn^{\prime}}\right)$
$+\Omega\,\bar{{\triangle}}\Omega+(d-1)g^{mn}\partial_{m}\Omega\,\partial_{n}\Omega\,,$
where $\bar{{\triangle}}$ indicates $(D-d)$-dimensional Laplace operator with respect to the metric $g_{mn}$. We will try to obtain a restriction on the values of the acceleration using the strong energy condition in $D$-dimensions. It gives
$0\leq R_{tt}$ $=$ $-\left(d-1\right)\left[a^{-1}\ddot{a}+\left(a\,\Omega\right)^{-1}\dot{a}\,\dot{\Omega}-\Omega^{-2}\,\dot{\Omega}^{2}+\Omega^{-1}\,\ddot{\Omega}\right]$ (8)
$-\frac{1}{2}\left(g^{mn}\ddot{g}_{mn}+\dot{g}^{mn}\dot{g}_{mn}-\Omega^{-1}\,g^{mn}\dot{g}_{mn}\,\dot{\Omega}-\frac{1}{2}g^{mm^{\prime}}\,g^{nn^{\prime}}\,\dot{g}_{nm^{\prime}}\dot{g}_{mn^{\prime}}\right)$
$+\Omega\,\bar{{\triangle}}\Omega+(d-1)g^{mn}\partial_{m}\Omega\,\partial_{n}\Omega\,,$
where we have used the $D$-dimensional Einstein equations.
With multiplying Eq. (8) by $\Omega^{d-2}$, integrating over the compact manifold and using a residual diffeomorphism (5), we have
$(d-1)a^{-1}\ddot{a}\leq\int d^{D-d}y\Omega^{d-2}\sqrt{\det g_{mn}}\left[\Omega^{-1}\left\{\ddot{\Omega}+(d-1)H\dot{\Omega}\right\}+2(d-2)\Omega^{-2}\dot{\Omega}^{2}\right.$
$\left.~{}~{}~{}~{}~{}~{}+\frac{1}{4}\left(g^{mn}\dot{g}_{mn}\right)^{2}+\frac{1}{2}\left(\Omega^{-1}g^{mn}\dot{g}_{mn}\dot{\Omega}+\frac{1}{2}g^{mm^{\prime}}g^{nn^{\prime}}\dot{g}_{nm^{\prime}}\dot{g}_{mn^{\prime}}\right)\right],$ (9)
where we have used the fact that the integral of the Laplacian of the warp factor over the whole manifold is zero, because the manifold has no boundary and the warp factor is non-singular.
We can obtain a lower bound on the second derivative of the warp factor which is related to the acceleration of the scalar fields by using the $D$-dimensional Einstein equations. Here we will discuss the $d$-dimensional effective action to relate this restriction to a bound on the potential.
Since the inequality (9) gives a bound on the value of the scalar field acceleration only in the direction of the overall volume change, other deformations of the manifold are not important. We will define a scalar mode $\tau$ as a basis, which is the overall volume modulus while other fields are chosen to be orthogonal to it. Now it is necessary for us to find the $d$-dimensional effective action in order to relate the bound on the acceleration to a bound on the derivative of the potential. As was noted above, only overall scaling modulus $\tau$ will be analyzed. Separating the modulus $\tau$ from the others, we find
$ds^{2}$ $=$ ${\rm e}^{-2\sqrt{\frac{D-d}{(D-2)(d-2)}}\,\tau}\,\tilde{\Omega}^{2}(\Phi,y)\left[-dt^{2}+a^{2}(t)\delta_{\alpha\beta}dx^{\alpha}dx^{\beta}\right]$ (10)
$+{\rm e}^{2\sqrt{\frac{d-2}{(D-d)(D-2)}}\,\tau}\,\tilde{g}_{mn}(\Phi,y)dy^{m}dy^{n}\,.$
Here, the tilde denotes that the overall modulus contribution was singled out. The relevant part of the $d$-dimensional effective action is
$S=\int d^{d}x\sqrt{-{\rm det}\,g_{d}}\left[\frac{1}{2\kappa_{d}^{2}}R_{d}+\frac{1}{2\kappa_{d}^{2}}\dot{\tau}^{2}-V(\tau,\Phi)+\cdots\right],$ (11)
where $\kappa_{d}$ is the $d$-dimensional gravitational coupling constant and $\Phi$ denotes for all other fields. The $d$-dimensional Einstein equations and the equation of motion for the scalar field $\tau$ are
$R_{tt}=-(d-1)a^{-1}\ddot{a}=\dot{\tau}^{2}-\frac{2\kappa_{d}^{2}}{d-2}V\,,$ (12a)
$R_{ij}=\left[a^{-1}\ddot{a}+(d-2)H^{2}\right]g_{ij}=\frac{2\kappa_{d}^{2}}{d-2}g_{ij}V\,,$ (12b)
$\ddot{\tau}+(d-1)H\dot{\tau}+\kappa_{d}^{2}\partial_{\tau}V=0\,.$ (12c)
From the Einstein equations (12), we have
$H^{2}=\frac{1}{(d-1)(d-2)}\left(\dot{\tau}^{2}+2\kappa_{d}^{2}V\right)\,,~{}~{}~{}a^{-1}\ddot{a}=H^{2}-(d-2)^{-1}\dot{\tau}^{2}\,.$ (13)
Then, we obtain
$a^{-1}\ddot{a}=\frac{1}{(d-1)(d-2)}\left[-(d-2)\dot{\tau}^{2}+2\kappa_{d}^{2}V\right].$ (14)
The left hand side of Eq. (9) becomes
$(d-1)a^{-1}\ddot{a}=-\dot{\tau}^{2}+\frac{2\kappa_{d}^{2}}{d-2}V\,.$ (15)
If we use the scalar field equation (12c) and the definition of modulus $\tau$ , we have
$\Omega(t,y)$ $=$ $\exp\left[-\sqrt{\frac{D-d}{(D-2)(d-2)}}\,\tau\right]\tilde{\Omega}(\Phi,y)\,,$ (16a)
$g_{mn}(t,y)$ $=$ $\exp\left[\sqrt{\frac{d-2}{(D-d)(D-2)}}\,\tau\right]\tilde{g}_{mn}(\Phi,y)\,.$ (16b)
The right hand side of Eq. (9) can be rewritten as
$\left[\frac{-d^{3}+d^{2}(D-1)+6D(d-2)+4}{4(D-2)(d-2)}\right]\dot{\tau}^{2}+\sqrt{\frac{D-d}{(D-2)(d-2)}}\kappa_{d}^{2}\partial_{\tau}V\,.$ (17)
Combining Eqs. (15) and (17) with the SEC bound in Eq. (9), we obtain the following bound on the gradient of the potential,
$-\frac{\partial_{\tau}V}{V}\geq 2\sqrt{\frac{D-2}{(D-d)(d-2)}}-\alpha\frac{\dot{\tau}^{2}}{V}\,,$ (18)
where $\alpha$ is defined by
$\alpha=\frac{1}{4\kappa_{d}^{2}}\sqrt{\frac{D-d}{(D-2)(d-2)}}\left[-5d^{2}+d\left(5D+12\right)-11D-2\right].$ (19)
If the scalar field $\tau$ satisfies the slow-roll condition
$\frac{\dot{\tau}^{2}}{V}\ll 1\,,$ (20)
the bound in Eq. (18) reduces to the well-known swampland criteria Obied:2018sgi
$-\frac{\partial_{\tau}V}{V}\geq 2\sqrt{\frac{D-2}{(D-d)(d-2)}}=\lambda_{\rm SEC}\,.$ (21)
When the velocity of the scalar field is not negligible, we have to use Eq. (18) that we call dynamical de Sitter conjecture.
III Application to Quintessence
In this section, we discuss an application of dynamical de Sitter conjecture in Eq. (18) obtained in the last section.
As an example in which dynamics of the scalar field is essential, here we focus on the quintessence model for inflation. In such a case, Eqs. (12c) and (13) can be solved exactly which the $d$-dimensional exponential potential Abbott:1984fp ; Lucchin:1984yf ; Lucchin:1985wy ; Lyth:1991bc
$V=V_{0}\,\exp\left(-\lambda\tau\right)\,,$ (22)
with $V_{0}$ and $\lambda>0$ arbitrary constants. It is easy to verify that Eqs. (12c) and (13) are satisfied by
$\tau(t)=\frac{1}{\lambda}\ln\left(\tau_{0}t^{2}\right)\,,~{}~{}~{}~{}~{}\tau_{0}=\frac{2\kappa_{d}^{2}\,V_{0}\,\varepsilon^{2}}{(d-2)(d-1-\varepsilon)}\,,$ (23)
and
$a\propto t^{1/\varepsilon}\,,~{}~{}~{}~{}~{}H=\frac{1}{\varepsilon t}\,,$ (24)
where $\varepsilon$ is the positive dimensionless quantity
$\varepsilon\equiv\frac{d-2}{4}\,\lambda^{2}\,.$ (25)
The extra term $\dot{\tau}^{2}/V$ becomes
$\frac{\dot{\tau}^{2}}{V}=\frac{4\tau_{0}}{\lambda^{2}\,V_{0}}=-\frac{2(d-2)\kappa_{d}^{2}\lambda^{2}}{4-2\lambda^{2}+d(\lambda^{2}-4)}\,.$ (26)
The value of the term $\frac{\dot{\tau}^{2}}{V}$ for each total dimension $D$ is listed in table 1 , where we set $d=4$ and $\kappa_{4}^{2}=1$ and $\lambda$ takes the smallest positive value in each dimension. One can see that as $D$ decreases, the kinetic term of the scalar field $\frac{\dot{\tau}^{2}}{V}$ cannot be neglected for any value of $V_{0}$
Table 1: We show the contribution of $\dot{\tau}^{2}/V$ for $d=4$ , $\kappa_{4}^{2}=1$ , when $\lambda$ is the smallest positive value. As the number of total dimension $D$ decreases, the term $\dot{\tau}^{2}/V$ becomes large.
$D$ 5 6 7 8 9 10 11
$\dot{\tau}^{2}/V$ 0.945619 0.4 0.253046 0.185636 0.147017 0.121975 0.104402
The potential of the scalar field depends on time
$V(t)=\bar{V_{0}}\,t^{-2}\,,$ (27)
where $\bar{V_{0}}$ denotes
$\bar{V_{0}}\equiv\frac{V_{0}}{\tau_{0}}=\frac{8\left(d-1-\frac{d-2}{4}\lambda^{2}\right)}{(d-2)\kappa_{d}^{2}\,\lambda^{4}}\,.$ (28)
Upon setting
$b=2\sqrt{\frac{D-2}{(D-d)(d-2)}}\,,$ (29)
we can find that the swampland criteria (18) gives
$\lambda\geq b-\frac{2\alpha\,\kappa_{d}^{2}\,\lambda^{2}}{\frac{4(d-1)}{d-2}-\lambda^{2}}\,.$ (30)
Then, the slope of the scalar potential $\lambda$ has to obey
$\lambda^{3}-(2\alpha\,\kappa_{d}^{2}+b)\lambda^{2}-\frac{4(d-1)}{d-2}\kappa_{d}^{2}\lambda+\frac{4(d-1)}{d-2}b\leq 0\,.$ (31)
In a setup of string theory ($D=10,d=4$), $\lambda$ should satisfy
$\lambda\leq-0.875001\,,~{}~{}~{}~{}0.587274\leq\lambda\leq 19.0671\,.$ (32)
(a)                                            (b)                                                (c)
Figure 1: The allowed region of the parameter $\lambda$ for $d=4$ is depicted in the case of $D=11$ (a), $D=10$ (b), and $D=5$ (c). One can recognize that the lower bound of $\lambda$ increases as the number of total dimension $D$ decreases.
For the eleven-dimensional supergravity ($D=11,d=4$), we find
$-0.797148\geq\lambda\,,~{}~{}~{}~{}0.54559\leq\lambda\leq 22.1224\,.$ (33)
We also consider the five-dimensional supergravity, the the slope of the scalar potential $\lambda$ should satisfy
$-1.99667\geq\lambda\,,~{}~{}~{}~{}~{}1.38786\leq\lambda\leq 5.30366\,.$ (34)
In the cases of exponential potentials that we investigated, an upper bound, $\lambda<\sqrt{3}$, was found from the minimal phenomenological requirements of a past epoch of radiation domination and acceleration today, while the simplest string theory realizations of the exponential potential in a flat universe have $\lambda>\sqrt{2}$ Agrawal:2018own ; Akrami:2018ylq ; Raveri:2018ddi ; Schoneberg:2023lun ; Andriot:2024jsh . On the other hand, there is constraint of $\lambda<0.85$ (95% CL) when using all currently available data such as BAO and Pantheon, and $\lambda<0.60$ (95% CL) when adding a prior on $H_{0}$ from Schoneberg:2023lun . However, surprisingly we find constraints to previous works Obied:2018sgi , leading to a constraint of $\lambda>0.6$ in the string theory. Thus, there is a mild tension on the de Sitter swampland criterion,
IV Summary and discussions
In this paper, we have discussed the de Sitter conjecture criteria when the time derivative of the scalar field cannot be neglected. When considering the quintessence model, this term is no longer negligible. In such cases, using the original de Sitter conjecture criteria Obied:2018sgi which is derived by completely ignoring the time derivative of the scalar field to discuss the validity of the quintessence model may no longer be entirely appropriate.
We have focused on the quintessence model with an exponential potential of the scalar field and modified the criteria of the de Sitter conjecture by considering the new contributions emerging from the time derivative of the scalar field. We have found that, particularly when the higher-dimensional spacetime is 5 or 6 dimensions, the contributions generated by the time derivative terms cannot be ignored. If we consider quintessence models, which always have the potential dominating over the kinetic term, we will encounter the fine-tuning and coincidence problems as the cosmological constant, lacking the basin of attraction of tracker models. Hence, it is necessary for us to treat a contribution of the kinetic term of scalar field for a successful model. Moreover, our criteria of de Sitter conjecture differ in the allowed region of potential slopes compared to the original de Sitter conjecture.
This offers the potential to resolve discrepancies between supernova and galaxy observations and four-dimensional dark energy models. Of course, our model only deals with a gravity and scalar field system, effectively ignoring the presence of other matter fields and radiation. Therefore, we cannot definitively assert that our results will immediately bridge the gap between observations and theory. However, by discussing the novel contributions proposed in this paper within a complete dynamical system, we may take a step closer to resolving these issues.
The dS conjecture Obied:2018sgi cannot hold because of of the problem with the Higgs potential that we discovered in Denef:2018etk . This has not yet been resolved (see also Choi:2018rze ). Our new criteria also fall victim to the Higgs potential problem.
In our paper, the strong energy condition is assumed in the higher-dimensional theory. We assume that there is no Higgs potential as described by Denef:2018etk in the higher-dimensional theory. If there is a Higgs potential in original higher-dimensional theory, we have to consider the equation of motion of the higher-dimensional Higgs field in the process of deriving equation (18) in our paper. However, if the Higgs potential of Denef:2018etk appears ”effectively” in the four-dimensional effective theory, the new criteria will not solve the Higgs problem. Even if the de Sitter criteria have a kinetic term of scalar field, the Higgs potential has a much larger contribution than the quintessence potential. The strong energy condition imposed in the paper will be violated at least at the extreme value of the Higgs potential. Even in this case, it would be ideal to choose the ansatz of the metric well and consider a compactification in which the Higgs and quintessence potentials are described by Denef:2018etk . Furthermore, if we impose a strong energy condition in the four-dimensional effective theory, it will give a constraint on $\lambda$. Regarding the second derivative of the potential of a scalar field, which is treated in Ooguri:2018wrx , it cannot be obtained simply by combining field equations and energy conditions. The dS criteria involving the second derivative are discussed using a simpler derivation of the asymptotic form and point out a potential loophole Hebecker:2018vxz .
Acknowledgements.
We thank Arthur Hebecker for valuable comments. The work of MN is supported in part by JSPS KAKENHI [Grants No. JP22H01221] and the WPI program “Sustainability with Knotted Chiral Meta Matter (SKCM2)” at Hiroshima University (KN and MN). The work of K. U. is supported by Grants-in-Aid from the Scientific Research Fund of the Japan Society for the Promotion of Science, under Contract No. 16K05364 and by the Grant “Fujyukai” from Iwanami Shoten, Publishers.